OpenStudy (anonymous):

\[\sqrt[4]y-7=2\]=17? i dont think i typed this right but check aND see. thanks.

OpenStudy (anonymous):

if this is \[\sqrt[4]{y-7}=2\] then \[y-7=2^4=16\] and so \[y=23\]